The B767-322ER aircraft has 88 more seats than the B767-333ER aircraft. The B767-332ER has 32 fewer seats than the B767-333ER aircraft. If their total number of seats is 332, find the number of seats for B767-322ER aircraft.
Answer #1

Let B767 -333ER has x seats .

Then B767 -322ER has (x +88) seats .

B767 - 332 ER has (x -32) seats .

Now by the problem ,

x + (x +88)+(x -32) = 332

x + x + x = 332 +32 -88

3x = 364 -88

3x = 276

x = 276/3

x = 92

Then B767-333ER has 92 seats .

B767-322ER has (92 +88) = 180 seats .

B767-322ER has 180 seats .